# SpeeDe3DGS: Speedy Deformable 3D Gaussian Splatting with Temporal Pruning and Motion Grouping

### v3

@@ Abstract
Dynamic extensions of 3D Gaussian Splatting (3DGS) achieve high-quality reconstructions through neural motion fields, but per-Gaussian neural inference makes these models computationally expensive. Building on DeformableGS, we introduce Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS), which bridges this efficiency–fidelity gap through three complementary modules: Temporal Sensitivity Pruning (TSP) removes low-impact Gaussians via temporally aggregated sensitivity analysis, Temporal Sensitivity Sampling (TSS) perturbs timestamps to suppress floaters and improve temporal coherence, and GroupFlow distills the learned deformation field into shared SE(3) transformations for efficient groupwise motion. On the 50 dynamic scenes in MonoDyGauBench, integrating TSP and TSS into DeformableGS accelerates rendering by 6.78 on average while maintaining neural-field fidelity and using 10 fewer primitives. Adding GroupFlow culminates in 13.71 faster rendering and 2.53 shorter training, surpassing all baselines in speed while preserving superior image quality.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) [mildenhall2021nerf] have transformed novel-view synthesis by representing static scenes as continuous volumetric functions optimized through differentiable rendering. While capable of photorealistic reconstruction, NeRFs rely on dense per-ray MLP inference, resulting in slow training and rendering. 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] replaces volumetric integration with differentiable rasterization of point-based Gaussian primitives, achieving real-time rendering with high fidelity and rapid convergence. This efficiency has made 3DGS a foundation for modern approaches to 3D reconstruction.
Dynamic Gaussian Splatting extends this framework to time-varying scenes by coupling each Gaussian with a learned motion field. Methods such as DeformableGS [yang2023deformable3dgs] and 4DGS [Wu2024CVPR4dgs] model temporal variation using neural deformation fields or spatiotemporal grids, enabling detailed dynamic reconstructions. However, these approaches require per-Gaussian neural inference at every frame, leading to substantial computational cost. Monocular Dynamic Gaussian Splatting Benchmark (MonoDyGauBench) [liang2025monocular] presents the first systematic comparison across dynamic 3DGS variants, revealing that neural motion fields achieve the highest and most stable reconstruction quality but render several times slower than analytic or non-neural representations.
@@ 2 Related Work
Neural Radiance Fields (NeRF) [mildenhall2021nerf] represent scenes as continuous volumetric functions optimized via differentiable volume rendering. Dynamic extensions incorporate time either through explicit temporal modeling [du2021nerflow, gao2021dnvs, wang2021dctnerf] or by learning deformation fields that map a canonical space to time-varying observations [pumarola2020dnerf, park2021nerfies, park2021hypernerf]. Grid-based and factorized encodings [guo2022neural, liu2022devrf, peng2023mlpmap, fang2022tineuvox, fridovich2023k, cao2023hexplane, gao2023strivec, mihajlovic2024resfields, wang2023msth] improve training and inference efficiency, but rendering remains slow due to per-ray MLP evaluation.
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] replaces volumetric integration with differentiable rasterization of Gaussian primitives, enabling real-time rendering for static scenes. Dynamic extensions incorporate explicit or implicit motion models to represent time-varying geometry. Dy3DGS [luiten2023dynamic3dgs] fits 6-DoF trajectories to individual Gaussians, while DeformableGS [yang2023deformable3dgs] and 4DGS [Wu2024CVPR4dgs] learn neural deformation fields using MLPs or HexPlane grids [cao2023hexplane]. RTGS [yang2023ICLR4dgs] encodes temporal variation implicitly within a 4D Gaussian distribution. MonoDyGauBench [liang2025monocular] unifies these formulations and identifies five representative motion parameterizations – EffGS [katsumata2023efficient], STG [li2023spacetime], DeformableGS [yang2023deformable3dgs], 4DGS [Wu2024CVPR4dgs], and RTGS [yang2023ICLR4dgs] – each trading off reconstruction fidelity and efficiency through distinct models of motion.
@@ 2.1 Gaussian Pruning
Redundancy in Gaussian-based reconstructions has motivated extensive pruning research to improve efficiency and reduce memory overhead [fan2024lightgaussian, fang2024mini, niemeyer2024radsplat, HansonTuPUP3DGS, HansonSpeedy]. Most approaches either learn per-Gaussian pruning masks [zhang2024lp3dgslearningprune3d, lee2024compact] or compute heuristic importance scores to remove low-contribution Gaussians [niemeyer2024radsplat, fan2024lightgaussian, fang2024mini, lin2024rtgs, girish2023eagles, ali2024elmgs, ali2024trimming, liu2024compgs, papantonakis2024reducing]. Recent extensions adapt these strategies to dynamic scenes by incorporating temporal variation into pruning metrics [liu2024lgslightweight4dgaussian, javed2024temporallycompressed3dgaussian, zhang2024megamemoryefficient4dgaussian]. PUP 3D-GS [HansonTuPUP3DGS] introduced a Hessian-based sensitivity score to quantify each Gaussian’s contribution to reconstruction, while Speedy-Splat [HansonSpeedy] refined this for pruning during training, achieving over faster rendering with minimal fidelity loss. Building on these sensitivity-based principles, our approach generalizes gradient-based importance analysis to dynamic 3D Gaussian Splatting, enabling pruning that accounts for motion-dependent redundancy across time.
@@ 2.2 Motion Analysis and Group-Based Modeling
Another strategy for compressing dynamic scenes is to reduce temporal complexity by decomposing motion into static and dynamic regions. Since most real-world scenes are largely static, several methods identify dynamic regions via motion cues [sabourgoli2024spotlesssplats, wang2025degauss, Zhou2024HUGS, xu2024das3r, wu2025swift4d] or assign Gaussians using self-supervised or segmentation-based clustering [wang2025degauss, xu2024das3r], significantly reducing computation by restricting 4D modeling to dynamic areas.
However, static–dynamic decomposition alone is insufficient for complex non-rigid motion. Group-based motion representations address this by jointly modeling correlated motion across sets of Gaussians. SC-GS [huang2023scgs] and SPGaussian [wan2024superpoint] apply control-point-based linear blend skinning with MLPs, RigGS [yao2025riggs] extracts sparse skeletons but retains a residual MLP for fine motion, and DynMF [kratimenos2024dynmf] learns optimizable polynomial motion bases. While effective, these methods still require costly neural inference. In contrast, our approach distills neural motion into grouped SE(3) transformations, clustering Gaussians by trajectory similarity to share rigid motion and reduce per-Gaussian inference while maintaining high fidelity.
@@ 3 Background
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] represents scenes as parametric, point-based models composed of 3D Gaussians. Given a set of ground truth training images , the scene is initialized using Structure from Motion (SfM) to produce image pose estimates and a sparse point cloud that serves as the initial means for the 3D Gaussians. Each image is paired with its corresponding pose in and used to optimize the scene.
@@
where is the projection of onto image space and is the inverse of the 2D covariance computed via the EWA Splatting approximation [zwicker2002ewa] of the perspectively projected 3D Gaussian. The model is optimized via stochastic gradient descent on an image reconstruction loss
@@ 3.2 Dynamic Gaussian Splatting
Recent dynamic Gaussian Splatting methods differ in how they parameterize , balancing motion expressiveness against computational efficiency. MonoDyGauBench [liang2025monocular] categorizes representative formulations into five major types:
@@
EffGS [katsumata2023efficient] models per-Gaussian motion using second-order Fourier and first-order polynomial bases.
@@
SpaceTimeGaussians (STG) [li2023spacetime] combines radial basis and polynomial bases to represent locally supported motion trajectories.
@@
DeformableGS [yang2023deformable3dgs] learns a shared MLP deformation field conditioned on spatial position and time.
@@
4D Gaussian Splatting (4DGS) [Wu2024CVPR4dgs] replaces the MLP with a factorized HexPlane grid [cao2023hexplane] for more efficient spatiotemporal interpolation.
@@
RTGS [yang2023ICLR4dgs] is the only approach without an explicit motion model, encoding temporal variation implicitly within a 4D Gaussian distribution.
@@ 4.1 Temporal Sensitivity Pruning
Recent studies have shown that 3D Gaussian Splatting models are heavily over-parameterized – comparable visual fidelity can often be achieved with far fewer primitives [HansonTuPUP3DGS]. Because the computational cost of deformable 3DGS scales linearly with the number of Gaussians, reducing redundant primitives is a direct path to accelerating both training and rendering. Building on the gradient-based pruning method of Speedy-Splat [HansonSpeedy], which achieves state-of-the-art efficiency in static 3DGS, we extend this concept to dynamic reconstruction. Our formulation introduces a temporal sensitivity pruning score that quantifies each Gaussian’s cumulative influence on reconstruction across both space and time.
@@
Here, is the set of training poses and timestamps , and is the rendered image for the deformed Gaussians at time and pose . is the 2D projection of Gaussian as defined in Equation 2, and is the image-space gradient readily available from the renderer’s backward pass. Since deformation parameters vary across timesteps, these gradients inherently encode temporal motion coupling, making sensitive not only to static appearance but also dynamic contribution.
Our approach – Temporal Sensitivity Pruning (TSP) – periodically removes Gaussians with low temporal sensitivity pruning scores, extending sensitivity-based pruning to dynamic 3DGS. By aggregating second-order sensitivities over both spatial and temporal domains, TSP identifies motion-dependent redundancies and eliminates them in a principled, temporally aware manner. This reduces training and inference cost while preserving visual fidelity by concentrating computation on the most informative Gaussians.
@@ 4.1.1 Temporal Sensitivity Sampling
Previous gradient-based analyses [HansonTuPUP3DGS, Jiang2023FisherRF, 11094666] have shown that gradient magnitudes can implicitly capture Gaussian sensitivity or stability. We extend this concept to dynamic 3DGS by interpreting low sensitivities as indicators of temporal inconsistency – Gaussians that appear stable under observed motion but drift when extrapolated to unseen deformation. To expose such instability, we introduce Temporal Sensitivity Sampling (TSS), a mechanism that probes the temporal dimension during sensitivity estimation.
@@
where is the deformation function, is standard normal noise, controls perturbation magnitude, is the frame interval, and is the total annealing period. In our TSS approach, we extend our TSP method by evaluating sensitivities at perturbed motion states defined by Equation 7. Since Equation 6 requires no ground-truth supervision, TSS can compute these sensitivities in a fully self-supervised manner by sampling any temporally jittered motion state.
This temporal perturbation reveals Gaussians that appear well-behaved on training frames but behave inconsistently under small motion shifts – the hallmark of floaters. Such primitives receive lower sensitivity scores and are pruned earlier, improving both spatial compactness and temporal coherence. Stronger perturbations early in training encourage temporal exploration and the removal of floaters during densification, while annealing the noise to zero in later iterations focuses optimization on precise reconstruction of the observed frames. This schedule balances temporal robustness with reconstruction accuracy, yielding stable and deterministic evaluations.
@@ 4.2.1 Flow Grouping via Motion Analysis
Next, we estimate a time-varying rigid transformation for each group to capture its motion across time. For each group and time , we estimate an SE(3) rigid transformation that maps this group from the canonical frame at time to time using Umeyama alignment [sorkine2007arap, umeyama1991align]. Specifically, we randomly sample a subset of means () from each group and estimate their rigid transformation via:
@@ 4.2.2 Flow Group Training and Inference
We set the shared flow of each group as learnable parameters. This proposed approach – GroupFlow – reduces the number of transformations predicted per timestep from (one per Gaussian) to (one per group) by enabling all Gaussians within a group to share the same motion parameters. We use groups to balance visual fidelity with total model size, as analyzed in Appendix A.3.
@@ 4.3 Speedy Deformable 3D Gaussian Splatting
We integrate our methods into DeformableGS [yang2023deformable3dgs] to form Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS), an efficient pipeline for dynamic scene reconstruction. During training, we apply Temporal Sensitivity Pruning (TSP) and Temporal Sensitivity Sampling (TSS) to progressively adapt model capacity to scene complexity. This two-phase strategy – soft pruning during densification and hard pruning afterward – removes redundant Gaussians and reduces deformation overhead, enabling real-time rendering while preserving the visual fidelity of neural motion models. After densification, we apply GroupFlow to further accelerate inference by distilling the learned deformation field into grouped SE(3) transformations. These shared groupwise motions replace per-Gaussian neural predictions, yielding rendering speeds that surpass prior dynamic Gaussian Splatting methods while maintaining high reconstruction quality and compact model size. We ablate the components of SpeeDe3DGS and compare them with other dynamic Gaussian Splatting approaches in Section 5.
@@ 5 Experiments
We evaluate our approach in two complementary settings: (i) our Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS) framework, which extends DeformableGS [yang2023deformable3dgs] for detailed ablation and analysis, and (ii) the Monocular Dynamic Gaussian Splatting Benchmark (MonoDyGauBench) [liang2025monocular], which standardizes comparisons across motion representations. Both settings integrate the complete SpeeDe3DGS pipeline – Temporal Sensitivity Pruning (TSP), Temporal Sensitivity Sampling (TSS), and GroupFlow – into dynamic Gaussian Splatting.
For consistent runtime comparison, all FPS metrics are measured on an NVIDIA RTX 3090. Unless otherwise specified, experiments follow each framework’s default hyperparameters and train for a total of 30,000 iterations. Starting at iteration 6,000, we apply TSP every 3,000 iterations, soft-pruning 60% of Gaussians at a time during densification and hard-pruning 30% after densficiation. The impact of different soft and hard pruning percentages is analyzed in Appendix A.4. Temporal perturbations for TSS are annealed using and , and GroupFlow is initialized after densification at iteration 15,000 with motion groups. The influence of group count on performance and model size is examined in Appendix A.3. This configuration enables detailed ablations within SpeeDe3DGS and ensures fair cross-method comparisons under MonoDyGauBench.

### v4

@@ Abstract
Dynamic extensions of 3D Gaussian Splatting (3DGS) achieve high-quality reconstructions through neural motion fields, but per-Gaussian neural inference makes these models computationally expensive. Building on DeformableGS, we introduce Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS), which bridges this efficiency-fidelity gap through three complementary modules: Temporal Sensitivity Pruning (TSP) removes low-impact Gaussians via temporally aggregated sensitivity analysis, Temporal Sensitivity Sampling (TSS) perturbs timestamps to suppress floaters and improve temporal coherence, and GroupFlow distills the learned deformation field into shared SE(3) transformations for efficient groupwise motion. On the 50 dynamic scenes in MonoDyGauBench, integrating TSP and TSS into DeformableGS accelerates rendering by 6.78 on average while maintaining neural-field fidelity and using 10 fewer primitives. Adding GroupFlow culminates in 13.71 faster rendering and 2.53 shorter training, surpassing all baselines in speed while preserving superior image quality.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) [33] have transformed novel-view synthesis by representing static scenes as continuous volumetric functions optimized through differentiable rendering. While capable of photorealistic reconstruction, NeRFs rely on dense per-ray MLP inference, resulting in slow training and rendering. 3D Gaussian Splatting (3DGS) [20] replaces volumetric integration with differentiable rasterization of point-based Gaussian primitives, achieving real-time rendering with high fidelity and rapid convergence. This efficiency has made 3DGS a foundation for modern approaches to 3D reconstruction.
Dynamic Gaussian Splatting extends this framework to time-varying scenes by coupling each Gaussian with a learned motion field. Methods such as DeformableGS [54] and 4DGS [49] model temporal variation using neural deformation fields or spatiotemporal grids, enabling detailed dynamic reconstructions. However, these approaches require per-Gaussian neural inference at every frame, leading to substantial computational cost. Monocular Dynamic Gaussian Splatting Benchmark (MonoDyGauBench) [25] presents the first systematic comparison across dynamic 3DGS variants, revealing that neural motion fields achieve the highest and most stable reconstruction quality but render several times slower than analytic or non-neural representations.
@@ 2 Related Work
Neural Radiance Fields (NeRF) [33] represent scenes as continuous volumetric functions optimized via differentiable volume rendering. Dynamic extensions incorporate time through explicit temporal modeling [4, 9, 44] or deformation fields that map a canonical space to time-varying observations [39, 36, 37]. Grid-based and factorized encodings [13, 29, 38, 7, 8, 3, 11, 32, 45] improve efficiency, but rendering remains slow due to per-ray MLP evaluation.
3D Gaussian Splatting (3DGS) [20] replaces volumetric integration with differentiable rasterization of Gaussian primitives, enabling real-time rendering for static scenes. Dynamic extensions incorporate explicit or implicit motion models to represent time-varying geometry. Dy3DGS [31] fits 6-DoF trajectories to individual Gaussians, while DeformableGS [54] and 4DGS [49] learn neural deformation fields using MLPs or HexPlane grids [3]. RTGS [53] encodes temporal variation implicitly within a 4D Gaussian distribution. MonoDyGauBench [25] unifies these formulations and identifies five representative motion parameterizations – EffGS [19], STG [24], DeformableGS [54], 4DGS [49], and RTGS [53] – each trading off reconstruction fidelity and efficiency through distinct models of motion.
@@ 2.1 Gaussian Pruning
Redundancy in Gaussian-based reconstructions has motivated extensive pruning research to improve efficiency and reduce memory overhead [5, 6, 34, 15, 14]. Most approaches either learn per-Gaussian pruning masks [56, 22] or compute heuristic importance scores to remove low-contribution Gaussians [34, 5, 6, 26, 12, 1, 2, 30, 35]. Recent extensions adapt these strategies to dynamic scenes by incorporating temporal variation into pruning metrics [28, 17, 55]. PUP 3D-GS [15] introduced a Hessian-based sensitivity score to quantify each Gaussian’s contribution to reconstruction, while Speedy-Splat [14] refined this for pruning during training, achieving over faster rendering with minimal fidelity loss. Building on these sensitivity-based principles, our approach generalizes gradient-based importance analysis to dynamic 3D Gaussian Splatting, enabling pruning that accounts for motion-dependent redundancy across time.
@@ 2.2 Motion Analysis and Group-Based Modeling
Another strategy for compressing dynamic scenes is to reduce temporal complexity by decomposing motion into static and dynamic regions. Since most real-world scenes are largely static, several methods identify dynamic regions via motion cues [40, 47, 57, 51, 50] or cluster Gaussians using self-supervised or segmentation-based methods [47, 51], restricting expensive 4D modeling to dynamic areas.
However, static–dynamic decomposition alone is insufficient for complex non-rigid motion. Group-based motion representations address this by jointly modeling correlated motion across sets of Gaussians. SC-GS [16], SPGaussian [43], and MoSca [23] model motion using control points with linear blend skinning or dual quaternion blending, often combined with MLP deformation. Methods such as DynMF [21], Gaussian-Flow [27], and Shape of Motion [46] instead learn low-dimensional motion bases or flow fields. While effective for improving visual quality, these methods require costly neural inference [16, 43], complex deformation computation [23], or guidance from 2D priors [23, 46]. In contrast, our approach distills neural motion into efficient grouped SE(3) transformations, clustering Gaussians by trajectory similarity to share rigid motion and reduce per-Gaussian inference while maintaining high fidelity.
@@ 3 Background
3D Gaussian Splatting (3DGS) [20] represents scenes as parametric, point-based models composed of 3D Gaussians. Given a set of ground truth training images , the scene is initialized using Structure from Motion (SfM) to produce image pose estimates and a sparse point cloud that serves as the initial means for the 3D Gaussians. Each image is paired with its corresponding pose in and used to optimize the scene.
@@
where is the projection of onto image space and is the inverse of the 2D covariance computed via the EWA Splatting approximation [58] of the perspectively projected 3D Gaussian. The model is optimized via stochastic gradient descent on an image reconstruction loss
@@ 3.2 Dynamic Gaussian Splatting
Recent dynamic Gaussian Splatting methods differ in how they parameterize , balancing motion expressiveness against computational efficiency. MonoDyGauBench [25] categorizes representative formulations into five major types:
@@
EffGS [19] models per-Gaussian motion using second-order Fourier and first-order polynomial bases.
@@
SpaceTimeGaussians (STG) [24] combines radial basis and polynomial bases to represent locally supported motion trajectories.
@@
DeformableGS [54] learns a shared MLP deformation field conditioned on spatial position and time.
@@
RTGS [53] is the only approach without an explicit motion model, encoding temporal variation implicitly within a 4D Gaussian distribution.
@@ 4.1 Temporal Sensitivity Pruning
Recent studies have shown that 3D Gaussian Splatting models are heavily over-parameterized – comparable visual fidelity can often be achieved with far fewer primitives [15]. Because the computational cost of deformable 3DGS scales linearly with the number of Gaussians, reducing redundant primitives is a direct path to accelerating both training and rendering. Building on the gradient-based pruning method of Speedy-Splat [14], which achieves state-of-the-art efficiency in static 3DGS, we extend this concept to dynamic reconstruction. Our formulation introduces a temporal sensitivity pruning score that quantifies each Gaussian’s cumulative influence on reconstruction across both space and time.
@@
Here, is the set of training poses and timestamps , and is the rendered image for the deformed Gaussians at time and pose . is the 2D projection of Gaussian as defined in Equation 2, and is the image-space gradient readily available from the renderer’s backward pass. Since deformation parameters vary across timesteps, these gradients inherently encode temporal motion coupling, making sensitive not only to static appearance but also to dynamic contribution.
Our Temporal Sensitivity Pruning (TSP) approach periodically removes Gaussians with low temporal sensitivity pruning scores, extending sensitivity-based pruning to dynamic 3DGS. By aggregating second-order sensitivities over both spatial and temporal domains, TSP identifies motion-dependent redundancies and eliminates them in a principled, temporally aware manner. This reduces training and inference cost while preserving visual fidelity by concentrating computation on the most informative Gaussians.
@@ 4.1.1 Temporal Sensitivity Sampling
Previous gradient-based analyses [15, 18, 48] have shown that gradient magnitudes can implicitly capture Gaussian sensitivity or stability. We extend this concept to dynamic 3DGS by interpreting low sensitivities as indicators of temporal inconsistency – Gaussians that appear stable under observed motion but drift when extrapolated to unseen deformation. To expose such instability, we introduce Temporal Sensitivity Sampling (TSS), a mechanism that probes the temporal dimension during sensitivity estimation.
@@
where is the deformation function, is standard normal noise, controls perturbation magnitude, is the frame interval, and is the annealing period. In our TSS approach, we extend our TSP method by evaluating sensitivities at perturbed motion states defined by Equation 7. Since Equation 6 requires no ground-truth supervision, TSS can compute these sensitivities in a fully self-supervised manner by sampling any temporally jittered motion state. Appendix A.4 ablates our choice of .
This temporal perturbation reveals Gaussians that appear well-behaved on training frames but behave inconsistently under small motion shifts – a hallmark of floaters. Such primitives receive lower sensitivity scores and are pruned, improving spatial compactness and temporal coherence. Stronger perturbations early in training encourage temporal exploration and the removal of floaters during densification, while annealing the noise to zero in later iterations focuses optimization on precise reconstruction of the observed frames. This schedule balances temporal robustness with reconstruction accuracy, yielding stable and deterministic evaluations.
@@ 4.2.1 Flow Grouping via Motion Analysis
Next, we estimate a time-varying rigid transformation for each group to capture its motion across time. For each group and time , we estimate an SE(3) rigid transformation that maps this group from the canonical frame at time to time using Umeyama alignment [41, 42]. Specifically, we randomly sample a subset of means () from each group and estimate their rigid transformation via:
@@ 4.2.2 Flow Group Training and Inference
where converts quaternions to rotation matrices and performs the inverse. We set the shared flow of each group as learnable parameters.
Our GroupFlow approach reduces the number of transformations predicted per timestep from (one per Gaussian) to (one per group) by enabling all Gaussians within a group to share the same motion parameters. In Appendix A.3, we ablate and select groups to balance visual fidelity with total model size.
@@ 4.3 Speedy Deformable 3D Gaussian Splatting
We integrate our methods into DeformableGS [54] to form Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS), an efficient pipeline for dynamic scene reconstruction. During training, we apply Temporal Sensitivity Pruning (TSP) and Temporal Sensitivity Sampling (TSS) to progressively adapt model capacity to scene complexity. This two-phase strategy – soft pruning during densification and hard pruning afterward – removes redundant Gaussians and reduces deformation overhead, enabling real-time rendering while preserving the visual fidelity of neural motion models. After densification, we apply GroupFlow to further accelerate inference by distilling the learned deformation field into grouped SE(3) transformations. These shared groupwise motions replace per-Gaussian neural predictions, yielding rendering speeds that surpass prior dynamic Gaussian Splatting methods while maintaining high reconstruction quality and compact model size. We ablate the components of SpeeDe3DGS and compare them with other dynamic Gaussian Splatting approaches in Section 5.
@@ 5 Experiments
We evaluate our approach in two complementary settings: (i) our Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS) framework, which extends DeformableGS [54] for detailed ablation and analysis, and (ii) the Monocular Dynamic Gaussian Splatting Benchmark (MonoDyGauBench) [25], which standardizes comparisons across motion representations. Both settings integrate the complete SpeeDe3DGS pipeline – Temporal Sensitivity Pruning (TSP), Temporal Sensitivity Sampling (TSS), and GroupFlow – into dynamic Gaussian Splatting.
For consistent runtime comparison, all FPS metrics are measured on an NVIDIA RTX 3090 GPU. Unless otherwise specified, experiments follow each framework’s default hyperparameters and train for a total of 30,000 iterations. Starting at iteration 6,000, we apply TSP every 3,000 iterations, soft pruning 60% of Gaussians at a time during densification and hard pruning 30% after densification. The impact of different soft and hard pruning percentages is analyzed in Appendix A.5. Temporal perturbations for TSS are annealed using and , and GroupFlow is initialized after densification at iteration 15,000 with motion groups. The influence of group count on performance and model size is examined in Appendix A.3. This configuration enables detailed ablations within SpeeDe3DGS and ensures fair cross-method comparisons under MonoDyGauBench.